OpenStudy (anonymous):

A data set consists of the following data points: (3,5), (5,8), (6,13) The slope of the best fit line is 2.5. Find the y-intercept of this line.

OpenStudy (mathmate):

There are standard formulae for the parameters of the regression line. One version is as follows. Assume y=mx+b, with n data points. m(slope) = (sum(x^2)sum(y)-sum(x)sum(xy))/(nsum(x^2)-(sum(x)^2) b (intercept)=(n*sum(xy)-(sum(x))(sum(y)))/(nsum(x^2)-(sum(x)^2) where sum(...) = summation over the data points. Using the given data, check that m=2.5, and subsequently calculate b. These formulae can be derived from first principles based on least squares and using basic calculus.